4. Compute the electric field at the origin due to the three point charges shown below. y (m) E3 Ey +2:00 AC 41 1.00 pC E1 0 -5.00 E2 Ex 3.50
Answer #1

Net horizontal electric field

Ex = k ( q1/ r1^2 + q2 / r2^2

Ex = 9*10^9* 10^-6* ( 2/ 5^2 + 1/3.5^2)

Ex = 1454.7 N/C

NEt vertical electric field

Ey = 9*10^9* 1.5*10^-6 / 4^2

Ey = 843.75 N/C

NEt electric field at origin

E^2 = Ex^2 + Ey^2

E = 1681.68

=========

Do comment in case any doubt, will reply for sure . Goodluck